Hi again,
My professor gave us this quadratic equation for homework:
a) x^2 - 2Rx + R^2 =0
He is asking us to solve for x and I am lost.
Thanks,
Louis

Found 2 solutions by stanbon, jake_6233:
Answer by stanbon(75887)

x^2 - 2Rx + R^2 =0
The left side is a "perfect square".
Rewrite as
(x-R)^2=0
Then x-R=0 twice
Then x=R with multiplicity two.
Cheers,
Stan H.

Answer by jake_6233(74)

x^2 - 2Rx + R^2 =0 This is a simple factoring job here
(x-R)(x-R)=0
when it comes to finding the zeros for this problem we have (x-R) twice so the answer is R with a multiplicity of 2 I don't know if this answer is beyond the scope of your class but this answer is complete. Please let me know if you need further clarification